OpenStudy (anonymous):

What is the rate of change for f(x)=-2cos4x-3 on the interval from x=pi/4 to x=pi/2?

OpenStudy (anonymous):

rate of change for f(x) is rate of change = [f(x2) - f(x1)] / [(x2 - x1)] here x2 = pi/2 and x1 = pi/4

OpenStudy (anonymous):

now f(x2) = -2cos(4.pi/2) - 3 = ................... and f(x1) = -2cos(4.pi/4) -3 = .................... can you do it now???

OpenStudy (anonymous):

Thanks, but my answer choices are: -16/pi, -pi/16, 2/pi, or pi/2

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):